OpenStudy (anonymous):

Solve. Please help me m^2-3=11m/2

OpenStudy (anonymous):

m=-.5 or m=6

OpenStudy (anonymous):

But how did you get that answer?

OpenStudy (anonymous):

Hello???

OpenStudy (anonymous):

m^2-3 = (11 m)/2 Subtract (11 m)/2 from both sides: m^2-(11 m)/2-3 = 0 Add 3 to both sides: m^2-(11 m)/2 = 3 Add 121/16 to both sides: m^2-(11 m)/2+121/16 = 169/16 Factor the left hand side: (m-11/4)^2 = 169/16 Take the square root of both sides: abs(m-11/4) = 13/4 Eliminate the absolute value: m-11/4 = -13/4 or m-11/4 = 13/4 Add 11/4 to both sides: m = -1/2 or m-11/4 = 13/4 Add 11/4 to both sides: m = -1/2 or m = 6

OpenStudy (anonymous):

Thanks & I see what I did wrong.